David bought a radio for $79.95. He paid $19.95 down payment and agreed to pay the balance plus an interest in three months. If the total interest he paid was $6, what is the rate of simple interest in three months?


1
Expert's answer
2021-11-16T11:06:51-0500

Cost of the radio=$79.95

Down payment=$19.95

Value of loan to purchase radio for 5 months

"=79.95-19.95=60"

Total interest paid for 3 months=6

Interest=loan value×time ×interest rate per annum

Time"=\frac{3}{12}" Years

"6=60\times\frac{3}{12}×interest rate"

"\frac{6×12}{60×3}=0.4"

Interest rate for 3months

"0.4×\frac{3}{12}"

=0.1 or 10%